Q.1 There is a country called Ecobliss. The economy consisted of 2 goods: A&B.This country is self suffiecient with no exports or imports. In 2019, the economy produced 20A and 50B. The prices (given here in the local currency rupiya) were 12 and 30 rupiya, respectively.
During the next year, the citizens of Ecobliss get struck by a deadly virus which induced extreme drowsiness in its victims. Productivity is hit hard due to this. In 2020 , the economy produce 15A at 15 rupiya each and 30B costing 32 rupiya each.
a.) Find the following information for 2019&2020.
Nominal GDP Real GDP(in 2019 prices)
GDP deflator (in 2019 prices)
Real GDP(in 2020 prices)
GDP deflator (in 2020 prices)
b) What is the rate of inflation between 2019 & 2020 using 2019 as a base year ?
c) What is the percentage change in the GDP deflator between 2019 &2020 using 2020 as the base year?
d) Why are your answers to b) and c) different
1
Expert's answer
2020-11-05T07:25:48-0500

Solution:

Q.1. a.). Nominal GDP

"Nominal \;GDP\;2019=sum \;of \;current\;year\;prices\;\times quantity \;of \;all\;the\;goods"

"Nominal\;GDP=(12\times 20)+ (30\times 50)"

"= 240+1500 = 1,740\;Rupiyas"


"Nominal\;GDP\;2020=(15\times 15)+ (32\times 30)"

"= 225+960 = 1,185\;Rupiyas"


"Real \;GDP=sum \;of \;the\;base\;year\;prices\;\times current\;year\;quantity \;of \;all\;the\;goods"


"Real \;GDP\;2019=(12\times 20)+ (30\times 50)"

"= 240+1500 = 1,740\;Rupiyas"


"Real \;GDP\;2020=(12\times 15)+ (30\times 30)"

"= 180+900 = 1,080\;Rupiyas"


"GDP\;deflector=(\frac{Nominal\;GDP}{Real\;GDP} ) \times 100"


"GDP\;deflector\;(in\; 2019\;prices) =100"

This is because 2019 is the base year and the deflector has to be equal to 100 without the need for calculations.


"GDP\;deflector\;(in\; 2019\;prices)=(\frac{1,185}{1,080} ) \times 100 = 109.7"


b.). The rate of inflation between 2019 and 2020 using 2019 as base year

Inflation rate using GDP reflector = the growth rate of the GDP reflector


"Inflation\; rate\;2019-2020=(\frac{109.7-100}{100} ) \times 100 = 9.7\%"


c.). The percentage change in the GDP deflator between 2019 &2020 using 2020 as the base year.

"GDP\;deflector\; 2019=(\frac{1,740}{1,740} ) \times 100 = 100"


"GDP\;deflector\; 2020=100\;since \;it\;is\;the\;base\;year"


"Percentage\;change\;in\;GDP\;deflector\;2019-2020=(\frac{100-100}{100} ) \times 100 = 0\%"


d.). The answers in b and c are different because of the change in the base year. Using the GDP deflector, changing the base year completely alters the inflation rate or the growth rate of GDP deflector.